# Krylov Complexity as a Probe for Chaos

### v2

@@ I Introduction
In an isolated quantum system defined by its Hamiltonian, along with specific boundary or initial conditions, it is natural to expect that the key indicators of its chaotic nature are encoded in the structure of its eigenstates and eigenvalues. Indeed, it is well known that the eigenvalues of chaotic Hamiltonians display distinct statistical features. More precisely, the level spacing of a maximally chaotic quantum system follows the Wigner-Dyson distribution, while that of integrable systems follows the Poisson distribution Bohigas et al. (1984).
@@
To investigate the nature of quantum chaos, several quantities have been introduced, including out-of-time- order correlators (OTOCs) Shenker and Stanford (2014a, b). Semiclassically, OTOCs exhibit exponential growth due to the butterfly effect, characterized by the Lyapunov exponent, which is conjectured to be bounded Maldacena et al. (2016). This bound is saturated in certain strongly interacting models with holographic descriptions, such as the Sachdev-Ye-Kitaev (SYK) model Kitaev ; Sachdev and Ye (1993); Maldacena and Stanford (2016). However, it is important to note that the exponential growth of OTOCs is not a universal characteristic of all chaotic systems Fine et al. (2014); Xu and Swingle (2019); Hashimoto et al. (2020).
@@
The present article aims to explore the chaotic nature of many-body quantum systems using Krylov complexity [^1] Throughout this paper, we will examine the Krylov complexity of states Balasubramanian et al. (2022); Alishahiha and Banerjee (2023); Caputa et al. (2024). , as defined in the context of the Lanczos algorithm. While this method has traditionally been employed for numerical computations Viswanath and Müller (1994), in recent years there has been increased interest in applying the Krylov method to the study of quantum chaos Parker et al. (2019) (for review see Nandy et al. (2024a))  [^2] For applications of Krylov method to entanglement generation and quantum metrology see e.g. Chu et al. (2024)..
@@
The authors of Parker et al. (2019) proposed a universal operator growth hypothesis that connects the asymptotic behaviour of Lanczos coefficients to the dynamics of the system in question. Specifically, for chaotic many-body quantum systems in dimensions greater than one and lacking symmetry, the Lanczos coefficients are expected to exhibit asymptotic linear growth. However, while the literature supports this linear growth in chaotic systems, it may not be a universal indicator of chaos; non-chaotic models can also display asymptotic linear growth Dymarsky and Smolkin (2021); Bhattacharjee et al. (2022a); Avdoshkin et al. (2024); Camargo et al. (2023); Vasli et al. (2024).
@@
The Krylov complexity saturates in many-body systems when the Lanczos coefficients vanish at certain steps. It is therefore natural to examine this saturation value to determine whether it can distinguish between chaotic and integrable models. Actually, in an explicit Ising model, it has been demonstrated that the saturation value of the Krylov complexity increases as one transitions from an integrable model to a chaotic one Rabinovici et al. (2022a, b). However, it is important to note that the saturation value of Krylov complexity also depends on the specific state for which it is computed. In fact, the saturation value for chaotic systems may or may not exceed that of integrable models Rabinovici et al. (2022a, b); Scialchi et al. (2024); Ballar Trigueros and Lin (2022); Español and Wisniacki (2023). We will return to this point later.
@@
Using the evolution of maximally entangled states in the Krylov basis, it was shown in Erdmenger et al. (2023) that the linear growth and saturation of Krylov complexity are generic features of many-body systems, regardless of whether they are chaotic or integrable. However, in chaotic systems, the Krylov complexity exhibits a peak before reaching its saturation value at late times Balasubramanian et al. (2022).
@@
The presence of this peak before saturation has recently been observed numerically in various models Huh et al. (2024); Camargo et al. (2024a); Nandy et al. (2024b); Bhattacharjee and Nandy (2024); Balasubramanian et al. (2024); Baggioli et al. (2024). It is important to note that the appearance of a peak prior to saturation is not solely determined by the system’s dynamics; it also depends on the initial state. For instance, it occurs for the initial state considered in Erdmenger et al. (2023). Nonetheless, it appears that the thermofield double state at infinite temperature provides a maximal universal bound on this peak Camargo et al. (2024a).
@@
In this paper, we aim to further explore how Krylov complexity can probe the chaotic nature of systems. To achieve this, we leverage the fact that Krylov complexity is essentially the expectation value of the number operator in the Krylov basis. Consequently, studying the time dependence of complexity reduces to analyzing the time evolution of this expectation value—an object we are more familiar with in the context of quantum thermalization (see e.g. Alishahiha (2023)).
@@
Inspired by the procedure that leads to the eigenstate thermalization hypothesis (ETH) Srednicki (1994); Deutsch (1991), we find that the growth and subsequent saturation of complexity at a time exponentially large in relation to the system size is a generic feature of Krylov complexity, regardless of whether the dynamics is chaotic or integrable. The specific aspect of complexity that is sensitive to the chaotic nature of the system is how it approaches its saturation value at late times. In chaotic systems, the saturation value of complexity reaches the infinite time average of the complexity at the saturation time, whereas in integrable models, it approaches this value from below over a relatively longer time.
@@ II Krylov basis vs energy eigenbasis
Starting with the same initial state $|\psi_{0}\rangle$ in the same quantum system with time-independent Hamiltonian $H$, the Krylov basis, $\{|n\rangle,n=0,1,2,\cdots,{\cal D}_{\psi}-1\}$ can be constructed as follows. The first element of the basis is identified with the initial state $|0\rangle=|\psi_{0}\rangle$ which we assume to be normalized. The other elements are then constructed, recursively, following the algorithm Viswanath and Müller (2008)
@@
This recursive procedure stops whenever $b_{n}$ vanishes which occurs for $n={\cal D}_{\psi}\leq{\cal D}$ which defines the dimension of Krylov space. Note that this algorithm produces an orthonormal and ordered basis together with coefficients $a_{n}$ and $b_{n}$ known as Lanczos coefficients Lanczos (1950).
@@
This can be argued by noting that the energy eigenstates can actually be organized into different blocks according to their symmetry charges. Therefore, selecting a particular initial state with a definite charge, its spreading through the Krylov basis nevertheless retains that same charge. The dimension of the Krylov space coincides with that of the corresponding symmetry block Herviou et al. (2021). As an immediate consequence of this fact, the summation in the equation (9) terminates at ${\cal D}_{\psi}$
@@
for $p,q=0,\cdots,n-1$. Substituting (15) into (12) and (14), one obtains an orthogonality relation for the function $g_{n}(E_{i})$ and a recursion relation, respectively, of the forms [^4] Equation (18) can be immediately recognized as the recursion relation of the continuant of a tridiagonal matrix. This is consistent with the fact that Hamiltonian has a tridiagonal form in Krylov basis Balasubramanian et al. (2023). We also observe that this recurrence relation actually resembles the three-term recurrence relation for Krawtchouk polynomials Miki et al. (2022). This relationship can, in turn, be utilized to derive an exact solution for $a_{n}$ and $b_{n}$.
@@ III Krylov complexity probes chaos
Krylov complexity is defined as the expectation value of the number operator in the Krylov basis Parker et al. (2019)
@@
[…]
@@
For finite many-body systems where the Lanczos coefficients vanish at the step $n={\cal D}_{\psi}$, the Krylov complexity saturates [^5] It is worth noting that for models where $b_{n}$ vanishes at $n={\cal D}_{\psi}$, one may have a symmetry so that $b_{n}=b_{{\cal D}_{\psi}-n}$. In this case, the vanishing of the Lanczos coefficients does not result in complexity saturation, instead, we get a periodic behaviour, which could be due to an underlying $SU(2)$ symmetry Caputa et al. (2022). Indeed, the exact solution of equation (18) based on Krawtchouk polynomials belongs to this category. Of course, in what follows, we will consider cases without such a symmetry, so that the complexity saturates at late times. . Therefore, at late times, we have
@@
where ${\cal C}_{\rm S}$ is the saturation value of complexity. To achieve such a behaviour, the time-dependent phase in equation (22) must sum to zero at late times, leading to a constant value equal to the late-time saturation value of the complexity. This constant should be contrasted with the infinite time average. Actually, from (22) one finds
@@
From equation (22) and following the observation that complexity saturation begins at exponentially large times, we can actually estimate the saturation value of complexity. Interestingly enough, unlike the previous general observation, the saturation value of complexity does actually depend on whether the system is chaotic or integrable, as we explore in what follows.
To proceed, we use the explicit form of the matrix elements of the number operator to rewrite the complexity (22) as follows [^6] To arrive at this result, we have used the fact that the expression for complexity obtained by plugging (26) into (22), is manifestly symmetric under the exchange of $j$ and $k$.
${\cal C}(t)=-2\sum_{n=0}^{{\cal D}_{\psi}-1}\frac{n}{\left(b_{n}!\right)^{2}}\sum_{j\neq k}^{{\cal D}_{\psi}}\sin^{2}(\frac{\omega_{jk}}{2}t)|c_{j}|^{2}|c_{k}|^{2}g_{n}(E_{j})g_{n}(E_{k})\,.$ (31)
@@
Since we are interested in the complexity at the large time limit, it is useful to arrange the energy eigenvalues in increasing order, $E_{1}<E_{2}<\cdots<E_{{\cal D}_{\psi}}$, and reformulate the summation in equation (31) as follows
@@
Since the eigenvalues are arranged in an increasing order, the smallest values of $s_{j+\ell}$, namely the ones with $\ell=1$ oscillate slower than those with higher $\ell$ in the late time limit. The contributions coming from these slow oscillating modes towards the late-time behaviour of complexity are thus determined by the level spacing of the system. It is well known that for integrable systems, the level spacing follows a Poisson distribution, while for chaotic systems, it conforms to the Wigner surmise [^7] In general for chaotic systems we have Wigner-Dyson distribution given by $P(s)=\mu_{0}s^{\delta}e^{-\mu_{1}s^{2}}$, for $\delta=1,2,3$ and ${\cal O}(1)$ constants $\mu_{0}$ and $\mu_{1}$ to normalize the distributions.
@@
In chaotic systems, the distribution of $s_{j+1}$ exhibits [^8] It is worth noting, however, $s_{j+l}$ appearing in (32) has the dimension of energy while the level spacing $s$ appearing in (33) is dimensionless. Of course, it is known that to find the universal level spacing an appropriate normalization is required Evnin and Piensuk (2019). a peak around a value which is of order one, and this behaviour extends to other $\ell$ as well. The $\sin^{2}$ factor oscillates between zero and one, and for large $t$, its average value approaches $\frac{1}{2}$. Thus, we can estimate the complexity ${\cal C}(t)$ at late times as
@@
$\dot{{\cal C}}(t)\approx-\sum_{n=0}^{{\cal D}_{\psi}-1}\frac{n}{\left(b_{n}!\right)^{2}}\sum_{j=1}^{{\cal D}_{\psi}-1}\sin(s_{j+1}t)\,s_{j+1}$ (38)
×$\hskip 88.2037pt\times\,|c_{j+1}|^{2}|c_{j}|^{2}g_{n}(E_{j+1})g_{n}(E_{j})\,.$
For integrable models, this expression approaches zero due to the $s_{\alpha}$ factor. In contrast, for chaotic models, due to the level spacing distribution and for certain initial states this could in general result in a finite value, which may be almost zero at a particular time $t_{m}$ identified with the saturation time. Moreover, it is also possible to show that the second derivative of this term is negative at that point, indicating that if such an extremum exists, it must be a maximum. Consequently, in this case, the complexity exhibits a peak before saturating at the value determined by the infinite time average of the complexity. Of course, it is important to note that, in general, depending on the initial state, the peak may not be present at all or even if it is present, it may not be pronounced enough.
@@
We note, however, that for certain initial states of a chaotic system, the complexity does exhibit a clear peck before saturation. To elaborate on this point further, we note that for any initial state $|\psi_{0}\rangle$ in a typical many-body system, one can define the canonical density matrix as follows

### v3

@@ I Introduction
In an isolated quantum system defined by its Hamiltonian, along with specific boundary or initial conditions, it is natural to expect that the key indicators of its chaotic nature are encoded in the structure of its eigenstates and eigenvalues. Indeed, it is well known that the eigenvalues of chaotic Hamiltonians display distinct statistical features. More precisely, the level spacing of a maximally chaotic quantum system follows the Wigner-Dyson distribution, while that of integrable systems follows the Poisson distribution Bohigas:1983er .
@@
To investigate the nature of quantum chaos, several quantities have been introduced, including out-of-time- order correlators (OTOCs) Shenker:2013yza ; Shenker:2013pqa . Semiclassically, OTOCs exhibit exponential growth due to the butterfly effect, characterized by the Lyapunov exponent, which is conjectured to be bounded Maldacena:2015waa . This bound is saturated in certain strongly interacting models with holographic descriptions, such as the Sachdev-Ye-Kitaev (SYK) model Kitaev ; Sachdev:1992fk ; Maldacena:2016hyu . However, it is important to note that the exponential growth of OTOCs is not a universal characteristic of all chaotic systems Fine_2014 ; Xu_2019 ; Hashimoto:2020xfr .
@@
The present article aims to explore the chaotic nature of many-body quantum systems using Krylov complexity [^1] Throughout this paper, we will examine the Krylov complexity of states Balasubramanian:2022tpr ; Alishahiha:2022anw ; Caputa:2024vrn . , as defined in the context of the Lanczos algorithm. While this method has traditionally been employed for numerical computations viswanath1994recursion , in recent years there has been increased interest in applying the Krylov method to the study of quantum chaos Parker:2018yvk (for review see Nandy:2024htc )  [^2] For applications of Krylov method to entanglement generation and quantum metrology see e.g. Chu:2024cah ..
@@
The authors of Parker:2018yvk proposed a universal operator growth hypothesis that connects the asymptotic behaviour of Lanczos coefficients to the dynamics of the system in question. Specifically, for chaotic many-body quantum systems in dimensions greater than one and lacking symmetry, the Lanczos coefficients are expected to exhibit asymptotic linear growth. However, while the literature supports this linear growth in chaotic systems, it may not be a universal indicator of chaos; non-chaotic models can also display asymptotic linear growth Dymarsky:2021bjq ; Bhattacharjee_2022 ; Avdoshkin:2022xuw ; Camargo:2022rnt ; Vasli:2023syq .
@@
The Krylov complexity saturates in many-body systems when the Lanczos coefficients vanish at certain steps. It is therefore natural to examine this saturation value to determine whether it can distinguish between chaotic and integrable models. Actually, in an explicit Ising model, it has been demonstrated that the saturation value of the Krylov complexity increases as one transitions from an integrable model to a chaotic one Rabinovici:2021qqt ; Rabinovici:2022beu . However, it is important to note that the saturation value of Krylov complexity also depends on the specific state for which it is computed. In fact, the saturation value for chaotic systems may or may not exceed that of integrable models Rabinovici:2021qqt ; Rabinovici:2022beu ; Scialchi:2023bmw ; Ballar_Trigueros_2022 ; Espa_ol_2023 . We will return to this point later.
@@
Using the evolution of maximally entangled states in the Krylov basis, it was shown in Erdmenger:2023wjg that the linear growth and saturation of Krylov complexity are generic features of many-body systems, regardless of whether they are chaotic or integrable. However, in chaotic systems, the Krylov complexity exhibits a peak before reaching its saturation value at late times Balasubramanian:2022tpr .
@@
The presence of this peak before saturation has recently been observed numerically in various models Huh:2023jxt ; Camargo:2024deu ; Nandy:2024wwv ; Bhattacharjee:2024yxj ; Balasubramanian:2024ghv ; Baggioli:2024wbz . It is important to note that the appearance of a peak prior to saturation is not solely determined by the system’s dynamics; it also depends on the initial state. For instance, it occurs for the initial state considered in Erdmenger:2023wjg . Nonetheless, it appears that the thermofield double state at infinite temperature provides a maximal universal bound on this peak Camargo:2024deu .
@@
In this paper, we aim to further explore how Krylov complexity can probe the chaotic nature of systems. To achieve this, we leverage the fact that Krylov complexity is essentially the expectation value of the number operator in the Krylov basis. Consequently, studying the time dependence of complexity reduces to analyzing the time evolution of this expectation value—an object we are more familiar with in the context of quantum thermalization (see e.g. Alishahiha:2022nhe ).
@@
Inspired by the procedure that leads to the eigenstate thermalization hypothesis (ETH) Srednicki:1994mfb ; Deutsch:1991 , we find that the growth and subsequent saturation of complexity at a time exponentially large in relation to the system size is a generic feature of Krylov complexity, regardless of whether the dynamics is chaotic or integrable. The specific aspect of complexity that is sensitive to the chaotic nature of the system is how it approaches its saturation value at late times. In chaotic systems, the saturation value of complexity reaches the infinite time average of the complexity at the saturation time, whereas in integrable models, it approaches this value from below over a relatively longer time.
@@ II Krylov basis vs energy eigenbasis
Starting with the same initial state $|\psi_{0}\rangle$ in the same quantum system with time-independent Hamiltonian $H$, the Krylov basis, $\{|n\rangle,n=0,1,2,\cdots,{\cal D}_{\psi}-1\}$ can be constructed as follows. The first element of the basis is identified with the initial state $|0\rangle=|\psi_{0}\rangle$ which we assume to be normalized. The other elements are then constructed, recursively, following the algorithm viswanath1994recursion
@@
This recursive procedure stops whenever $b_{n}$ vanishes which occurs for $n={\cal D}_{\psi}\leq{\cal D}$ which defines the dimension of Krylov space. Note that this algorithm produces an orthonormal and ordered basis together with coefficients $a_{n}$ and $b_{n}$ known as Lanczos coefficients Lanczos:1950zz .
@@
This can be argued by noting that the energy eigenstates can actually be organized into different blocks according to their symmetry charges. Therefore, selecting a particular initial state with a definite charge, its spreading through the Krylov basis nevertheless retains that same charge. The dimension of the Krylov space coincides with that of the corresponding symmetry block PhysRevB.103.134207 . As an immediate consequence of this fact, the summation in the equation (9) terminates at ${\cal D}_{\psi}$
@@
for $p,q=0,\cdots,n-1$. Substituting (15) into (12) and (14), one obtains an orthogonality relation for the function $g_{n}(E_{i})$ and a recursion relation, respectively, of the forms [^4] Equation (18) can be immediately recognized as the recursion relation of the continuant of a tridiagonal matrix. This is consistent with the fact that Hamiltonian has a tridiagonal form in Krylov basis Balasubramanian:2022dnj . We also observe that this recurrence relation actually resembles the three-term recurrence relation for Krawtchouk polynomials miki2022singleindexedexceptionalkrawtchoukpolynomials . This relationship can, in turn, be utilized to derive an exact solution for $a_{n}$ and $b_{n}$.
@@ III Krylov complexity probes chaos
Krylov complexity is defined as the expectation value of the number operator in the Krylov basis Parker:2018yvk
@@
$\overline{{\cal C}}=\lim_{T\rightarrow\infty}\frac{1}{T}\int_{0}^{T}\langle{\cal N}(t)\rangle\,dt={\rm Tr}(\rho_{\rm DE}{\cal N})\,.$ (24)
@@
For finite many-body systems where the Lanczos coefficients vanish at the step $n={\cal D}_{\psi}$, the Krylov complexity saturates [^5] It is worth noting that for models where $b_{n}$ vanishes at $n={\cal D}_{\psi}$, one may have a symmetry so that $b_{n}=b_{{\cal D}_{\psi}-n}$. In this case, the vanishing of the Lanczos coefficients does not result in complexity saturation, instead, we get a periodic behaviour, which could be due to an underlying $SU(2)$ symmetry Caputa:2021sib . Indeed, the exact solution of equation (18) based on Krawtchouk polynomials belongs to this category. Of course, in what follows, we will consider cases without such a symmetry, so that the complexity saturates at late times. . Therefore, at late times, we have
@@
where ${\cal C}_{\rm S}$ is the saturation value of complexity. This constant should be contrasted with the infinite time average of complexity $\overline{{\cal C}}$. Actually, to achieve such a behaviour, the time-dependent phase in equation (22) must sum to zero at late times, leading to a constant value equal to the late-time saturation value of the complexity. Indeed, from (22) one finds
@@
From equation (22) and following the observation that complexity saturation begins at exponentially large times, we can actually estimate the saturation value of complexity. To proceed, we use the explicit form of the matrix elements of the number operator to rewrite the complexity (22) as follows [^6] To arrive at this result, we have used the fact that the expression for complexity obtained by plugging (26) into (22), is manifestly symmetric under the exchange of $j$ and $k$.
${\cal C}(t)=\sum_{n=0}^{{\cal D}_{\psi}-1}\frac{-2n}{\left(b_{n}!\right)^{2}}\sum_{j\neq k}^{{\cal D}_{\psi}}\sin^{2}(\frac{\omega_{jk}}{2}t)|c_{j}|^{2}|c_{k}|^{2}g_{n}(E_{j})g_{n}(E_{k})\,.$ (31)
@@
Since we are interested in the complexity at the large time limit, it is useful to arrange the energy eigenvalues in increasing order, $E_{1}<E_{2}<\cdots<E_{{\cal D}_{\psi}}$, and reformulate the summation in equation (31) as follows777 We note that Eq. (32) naturally reproduces the expected behavior for energy eigenstates. In this case, the Krylov state expansion contains only one nonzero coefficient, say $c_{j_{0}}=1$, with all others zero. Since the complexity expression involves sums over products $|c_{j+\ell}|^{2}|c_{j}|^{2}$ with $\ell\geq 1$, each term in the sum vanishes, yielding ${\cal C}(t)=0$. This aligns perfectly with the physical intuition that an energy eigenstate remains static in Krylov space and hence has zero Krylov complexity.
@@
Since the eigenvalues are arranged in an increasing order, the smallest values of $s_{j+\ell}$, namely the ones with $\ell=1$ oscillate slower than those with higher $\ell$ in the late time limit. The contributions coming from these slow oscillating modes towards the late-time behaviour of complexity are thus determined by the level spacing of the system. It is well known that for integrable systems, the level spacing follows a Poisson distribution, while for chaotic systems, it conforms to the Wigner surmise888 In general, for chaotic systems, the level spacing distribution follows the Wigner-Dyson form $P(s)=\mu_{0}s^{\delta}e^{-\mu_{1}s^{2}}$, where $\delta=1,2,4$ corresponds to the Gaussian Orthogonal Ensemble (GOE), Gaussian Unitary Ensemble (GUE), and Gaussian Symplectic Ensemble (GSE), respectively. The constants $\mu_{0}$ and $\mu_{1}$ are ${\cal O}(1)$ normalization factors determined by the conditions $\int_{0}^{\infty}P(s)\,ds=1,\;\int_{0}^{\infty}sP(s)\,ds=1$, which ensure that $P(s)$ is a properly normalized probability distribution with unit mean spacing.
@@
In chaotic systems, the distribution of $s_{j+1}$ exhibits [^9] It is worth noting, however, $s_{j+l}$ appearing in (32) has the dimension of energy while the level spacing $s$ appearing in (33) is dimensionless. Of course, it is known that to find the universal level spacing an appropriate normalization is required Evnin:2018jbh . a peak around a value which is of order one, and this behaviour extends to other $\ell$ as well. The $\sin^{2}$ factor oscillates between zero and one, and for large $t$, its average value approaches $\frac{1}{2}$. Thus, we can estimate the complexity ${\cal C}(t)$ at late times as
@@
$\dot{{\cal C}}(t)\approx\sum_{j=1}^{{\cal D}_{\psi}-1}\sin(s_{j+1}t)\,s_{j+1}|c_{j+1}|^{2}|c_{j}|^{2}A(E_{j},E_{j+1})\,$, (38)
where $A(E_{j},E_{j+1})>0$ is
$A(E_{j},E_{j+1})=-\sum_{n=0}^{{\cal D}_{\psi}-1}\frac{n}{\left(b_{n}!\right)^{2}}g_{n}(E_{j+1})g_{n}(E_{j})\,.$ (39)
For integrable models, the expression (38) approaches zero due to the $s_{\alpha}$ factor. In contrast, for chaotic models, due to the level spacing distribution and for certain initial states this could in general result in a finite value, which may be almost zero at a particular time $t_{m}$ identified with the saturation time. Moreover, it is also possible to show that the second derivative of this term is negative at that point [^10] Note that for small but nonzero $s_{j+1}$—as is typical in chaotic systems—the extremum $t_{m}$ occurs near the point where $s_{j+1}t_{m}\sim\pi$. At this time, the derivative is negative because the $\cos(s_{j+1}t_{m})$ factor becomes negative, while all other terms in Eq. (38) remain positive. This implies a negative slope and supports the existence of a local maximum at $t=t_{m}$. , indicating that if such an extremum exists, it must be a maximum. Consequently, in this case, the complexity exhibits a peak before saturating at the value determined by the infinite time average of the complexity. Of course, it is important to note that, in general, depending on the initial state, the peak may not be present at all or even if it is present, it may not be pronounced enough.
@@
We note, however, that for certain initial states of a chaotic system, the complexity does exhibit a clear peak before saturation. To elaborate on this point further, we note that for any initial state $|\psi_{0}\rangle$ in a typical many-body system, one can define the canonical density matrix as follows